# Circuit-tuning: A Mechanistic Approach for Identifying Parameter Redundancy and Fine-tuning Neural Networks

### v2

@@ 5.1 The subject-verb disagreement Task
5.1.1 Task description and data preparation
The goal of the subject-verb disagreement task is to match a verb with a subject in an abnormal way. For example, “I is", “he are" and “the cows eats" are all expected results for this task. In each sentence, the token before the verb is called the END token. The automatic evaluation metric for this task is the logit difference between the flipped verb and the original verb at the END token. This is because the logit at the END token is directly used for predicting the verb.
The reason why we create this task as the start of our experiments is that we expect the model to acquire a new capability from scratch, and thus to study how the circuit evolves during fine-tuning. The task is simple and interesting for analyzing the learning dynamics with clear visualization.
Different from previous works [14, 34] that use template-based datasets, we collect real-world data from the Pile corpus [17] to ensure diversity and authenticity. We extract 60k sentences in the present tense in English and flip the forms of the verbs. We ensure the high quality of our dataset and believe it is meaningful for further research. For the details of our dataset, please refer to Appendix F.1.
5.1.2 Implementation
We use GPT2-small [47] in this task. We set the output of the attention and the MLP in each layer as upstream nodes, and the input of the query, key, value and the MLP in each layer as downstream nodes. During training, we use the logit difference discussed before as the metric $\mathcal{L}_{m}$. We follow Syed et al. [56] for path patching with mean ablation, with an improvement detailed in Appendix F.4.5. We sweep over a range of $N$ which is the number of edges to be saved during circuit discovery. Mini-batch SGD is used for optimization. We train the model under each setting for 3 epochs where the performance almost converges. For implementation details, please refer to Appendix F.2.
@@ 5.1.2 Implementation
Figure 1: Results in the subject-verb disagreement task.
5.1.3 Main results
@@
From Figure 1(a), we observe a flip in logit difference from negative to positive, which means the model adjusts its grammar to fit the data distribution of subject-verb disagreement. One noteworthy finding is that the model can generate abnormal texts in the past tense, such as “to be or not to be, that were a …”, which implies that the model really grasps the new grammar and applies it smartly, since there is no sentence in the present tense in our training data at all.
With the increasing number of top $N$ edges, the logit difference increases until $N$ approaches 1000. The number of tunable parameters is also saturated at this point. We believe that when $N=1000$, almost all the necessary parameters for this task are included, and the performance cannot be further improved, suggesting the minimality of the circuit and the existence of intrinsic dimensionality.
@@
We also track the perplexity (PPL) that is calculated over the test set with corpora in various scenarios. In Figure 1(b), we find that the PPL of full fine-tuning is high and fluctuates wildly during training, though the logit difference of it is higher. This implies that circuit-tuning provides better training stability as well as better preservation of the basic language modeling ability over full fine-tuning.
@@
Figure 2: The changes in indexes with various choices of top-N imply the existence of the node-level intrinsic dimensionality at $N=1000$ in a computational graph. Note that the parameter ratio varies and depends on the granularity of nodes. Theoretically more granular nodes result in lower parameter ratio because the location of the intrinsic nodes would be more precise.
5.1.4 Ablation study for the quality of the discovered circuit
To confirm the effectiveness of our method to find the required parameters, we (i) calculate the faithfulness and completeness [62] of the circuits and (ii) provide an ablation study to randomly unfreeze the nodes outside the circuit during training and check if the performance is influenced. Results of the above experiments in Appendix F.3 ensure that the circuits are discovered accurately. Besides, we observe an obvious turning point of the faithfulness and completeness in Figure 8 as $N$ increases. Also, we notice a similar and common turning point of logit difference, PPL and number of tunable parameters in Figure 2. All of the above observations show the shadow of intrinsic nodes.
5.1.5 Analyses on learning dynamics
(a) Illustration of the flip.
(b) Before fine-tuning
(c) After fine-tuning
Figure 3: The flip observed in Finding 2.
Finding 1: Interpretation of the circuits
Inspired by Wang et al. [62], we decompose the circuit into attention heads of different functions. Firstly we find out the heads that directly affect the output, namely the Subject Attribute Heads. Then we find out the heads responsible for the localization of the subject, namely the Subject Identification Heads. Finally, we find out heads that could affect the behaviors of the Subject Attribute Heads, namely the Collaborative Heads. For technique details and visualization of the heads, please refer to Appendix F.4.1.
Finding 2: The flip of the Subject Attribute Heads
The Subject Attribute Heads are responsible for matching the subject and the verb. We compute the dot product between the output $x$ of each attention head at the END token and $W_{U}(v_{flip})\!-\!W_{U}(v)$, the difference between the unembedding projections. Since the latter is fixed, we expect the projection of $x$ on the direction of it to be large, thus encouraging the probability difference between the two opposite verb forms (love v.s. loves, etc.). We visualize the heads before and after fine-tuning in Figure 3. Through comparison, we observe an obvious flip at head.10.9, implying the reversal in its function from subject-verb agreement to disagreement. Other heads (3.0, 6.0, 4.4, 11.8, etc) also see flips with varying degrees, serving as evidence for the self-adjustment of the functions inside the nodes. See Appendix F.4.1 for details.
Strengthened Edges Weakened Edges
Edge Dynamic change Edge Dynamic change
(a) circuit at 2000 steps
(b) circuit at 3000 steps
(c) circuit at 4000 steps
Figure 4: Evidence of Hebbian learning in the subject-verb disagreement task ($N=50$ edges shown).
Finding 3: The sharing of the Subject Identification Heads
The Subject Identification Heads attend heavily to the subject. One type of these heads attends to the END token (0.1, 0.3, etc), which is helpful to the cases like “they are"; the other type of heads (8.5, 10.5, etc) attends to the subject several tokens before, which is helpful to the cases like “the girl wearing a dress is". Both types of heads perform the same before and after fine-tuning. This implies that their functions are preserved and shared all the way down. Our finding is consistent with the study of intrinsic dimensionality in Aghajanyan et al. [1] that pre-training optimizes the description length without having direct access to the same tasks. The kind of consistency further confirms the feasibility of our theory in Section 3.
Finding 4: The interaction and collaboration inside the model
Interestingly, we observe that over the training process the nodes (attention heads) inside a circuit complete a task through a cooperative division of labor. Each head has its division of labor as discussed before. Besides, some heads directly affect the output, while others cooperate with them to affect the output indirectly. Several heads achieve a common goal through cooperation: they could affect a head only through combined effect, and adjust themselves during training. For details, please refer to Appendix F.4.1.
Finding 5: The evidence of Hebbian learning
We visualize the circuits during the fine-tuning process in Figure 4, and observe a phenomenon akin to Hebbian learning [10] that several edges are strengthened during training, just as the synapses between neurons can be strengthened after continuous stimulation. Also, several edges are weakened at the same time, as shown in Table 1. It is similar to the lateral inhibition among neurons [22] that an activated neuron can reduce the activity of its neighbors. We present the dynamics of the competitive learning and self-organization [25] inside a model, providing new discoveries to the generalization process on a new data distribution.
Due to limited space, all the above together with other interesting findings are detailed in F.4.
5.2 Application to Complex Tasks
Table 2: Experiment results of circuit-tuning on complex tasks.
Methods & Tasks Reasoning-based Reasoning-free Computation
Mathematics Logical Reasoning Gender De-biasing Reading Comprehension
@@ 5.2 Application to Complex Tasks
Acc@1 (%) F1 (%) Prejudice Risk $\downarrow$ Exact Match(%) Avg param ratio
Llama-3.2-1B-it 40.71 20.35 0.555 39.68 /
Llama-3.2-1B-it-full-tuning 46.47 26.89 0.533 36.73 1.00
Llama-3.2-1B-it-lora 44.58 22.51 0.530 34.30 1.79e-2
Llama-3.2-1B-it-circuit-tuning 45.56 27.06 0.312 41.78 7.65e-2
Llama-3.2-3B-it 70.36 42.71 0.641 54.12 /
Llama-3.2-3B-it-full-tuning 75.44 47.32 0.632 55.63 1.00
Llama-3.2-3B-it-lora 73.54 46.27 0.638 54.73 1.49e-2
Llama-3.2-3B-it-circuit-tuning 74.35 47.59 0.417 56.58 8.57e-2
@@
Llama-3.1-8B-it-full-tuning 83.76 49.53 0.640 59.60 1.00
Llama-3.1-8B-it-lora 80.21 47.64 0.643 58.97 1.03e-2
Llama-3.1-8B-it-circuit-tuning 82.97 50.54 0.420 60.04 9.37e-2
Task descriptions
We prepare two types of tasks based on whether reasoning is involved. For reasoning-based tasks, we focus on mathematics and logical reasoning with chain-of-thought (CoT) included. For reasoning-free tasks, we prepare a gender de-biasing task that requires the model to develop an unbiased perspective on genders, and a reading comprehension task that requires only the keywords as the answer. Task settings and evaluation metrics are detailed in Appendix G.1.
Figure 5: Comparison in gender debiasing results.
Figure 6: Evaluation on general/other capabilities.
Implementation
We apply our method to llama-3.2-1B/3B and llama-3.1-8B [11]. In addition, we compare it with full fine-tuning and LoRA. We set them as baselines for simple comparisons only to verify that our method is also able to locate the key parameters precisely, as discussed in Section 4.2.
For circuit-tuning, the graph settings of our method are the same as before in Section 5.1.2, except that each MLP layer is split into 64-dimensional heads. For gender de-biasing, We selectively add $\beta\cdot|\mathcal{L}_{m}|$ that acts as a regularization term in the loss $\mathcal{L}$ for explicit guidance and intervention during training (Appendix G.1.2). All other details like the choice of the patching metric $\mathcal{L}_{m}$, the number of edges $N$, and the hyper-parameters for circuit-tuning and baselines are detailed in Appendix G.2.
Main results Table 2 shows the results of each method, demonstrating that our method is capable of various scenarios. Hence it is possible to study the learning dynamics as in Section 5.1 on various scenarios. Amazingly, the task performance brought by circuit-tuning is better than LoRA, and even surpasses full fine-tuning in many tasks. Note that the computation of circuit-tuning depends on node granularity and is for reference only since our method is not dedicated to efficient fine-tuning.
Besides, with an interpretable intervention, the de-biasing result is clearly better than others, which is shown in Figure 6 and discussed in Appendix G.4. This suggests that circuit-tuning can be applied flexibly with task-specific conditions to achieve controllable and steerable training.
To check if other capabilities are affected during training, we evaluate on benchmarks involving general capabilities as well as reasoning, coding, and multilingual abilities. Results are shown in Figure 6 and detailed in Appendix G.3, indicating that circuit-tuning is more targeted at the key components in a model, ensuring its reliability and simplicity in interpretability research.

### v3

@@ 4.1 The subject-verb disagreement Task
To comprehensively investigate the the practicality of our method as well as the learning dynamics during fine-tuning, we first design a simple while interesting task called “subject-verb disagreement”. The goal of this task is to match a verb with a subject in an abnormal way. For example, “I is”, “he are” and “the cows eats” are all expected results for this task. In each sentence, the token before the verb is called the END token. The automatic evaluation metric for this task is the logit difference between the flipped verb $v_{flip}$ and the original verb $v$ at the END token:
$diff_{logit}=W_{U}(x)_{j}-W_{U}(x)_{i},\quad W_{U}\in\mathbb{R}^{D\times|\mathcal{V}|}$ (2)
where $x\in\mathbb{R}^{D}$ is the output of the last layer at the END token, $W_{U}$ is the unembedding matrix, and $i$ and $j$ are the indices of the original verb and the flipped verb in the vocabulary $\mathcal{V}$ of the language model. We use this metric because the logit at the END token is directly used for predicting the verb token. This task encourages the model to acquire a new capability based on the existed English grammar, and thus we can study in detail how the circuit evolves during fine-tuning.
4.2 Data preparation and implementation details
Different from previous works (Finlayson et al., 2021; Marks et al., 2024) that use template-based datasets, we collect real-world data from the Pile corpus (Gao et al., 2020) to ensure diversity and authenticity. We extract 60k sentences in the present tense in English and flip the forms of the verbs. We ensure the high quality of our dataset and believe it is meaningful for further research. For the details of our dataset and task definiitons, please refer to Appendix D.1.
We use GPT2-small Radford et al. (2019) in this task. We set the output of the attention and the MLP in each layer as upstream nodes, and the input of the query, key, value and the MLP in each layer as downstream nodes. During training, we use the logit difference discussed before as the metric $\mathcal{L}_{m}$ for the quantization of the final output during graph pruning. We follow Syed et al. (2023) for path patching with mean ablation, with an improvement detailed in Appendix D.4.5. We sweep over a range of $N$ which is the number of edges to be saved during graph pruning. Mini-batch SGD is used for optimization. We train the model under each setting for 3 epochs where the performance almost converges. More details can be found in Appendix D.2.
@@ 4.2 Data preparation and implementation details
Figure 2: The change in logit difference and PPL during training on the subject-verb disagreement task.
4.3 Main results
@@
From Figure 2(a), we observe a flip in logit difference from negative to positive, which means the model adjusts its grammar to fit the data distribution of subject-verb disagreement. One noteworthy finding is that the model can generate abnormal texts in the past tense, such as “to be or not to be, that were a …”, which implies that the model really grasps the new grammar and applies it smartly, since there is no sentence in the past tense in our training data at all. With the increasing number of top $N$ edges, the logit difference increases until $N$ approaches 1000. The number of tunable parameters is also saturated at this point. This implies that when $N=1000$, almost all the necessary parameters for this task are included, and the performance cannot be further improved, suggesting the minimality of the subgraph at this point.
@@
We also track the perplexity (PPL) calculated on the validation set. In Figure 2(b), we find that the PPL of full fine-tuning is high and fluctuates wildly during training, though the logit difference of it is higher. This implies that circuit-tuning provides better training stability as well as better preservation of the basic language modeling ability over full fine-tuning. This is because under identical experimental settings, circuit-tuning adjusts a smaller yet more critical subset of parameters compared to full fine-tuning. By minimally affecting the parameters corresponding to task-agnostic general capabilities, it exhibits greater stability during the learning and generalization process.
(a) llustration of the flip.
(b) Before fine-tuning
(c) After fine-tuning
Figure 3: Visualization of the flip of the Subject Attribute Heads. (a): A 2D illustration of the flip in the functionality of the Subject Attribute Heads. (b) (c): Heatmaps of the dot-product between the output of each attention head and $W_{U(v_{flip})}\!-\!W_{U(v)}$ before and after fine-tuning.
4.4 Analyses on learning dynamics
Finding 1: Interpretation of the circuits Inspired by Wang et al. (2022), we decompose the circuit into attention heads of different functions. Firstly we find out the heads that directly affect the output, namely the Subject Attribute Heads. Then we find out the heads responsible for the localization of the subject, namely the Subject Identification Heads. Finally, we find out heads that could affect the behaviors of the Subject Attribute Heads, namely the Collaborative Heads. For technique details and visualization of the heads, please refer to Appendix D.4.1.
Finding 2: The flip of the Subject Attribute Heads
The Subject Attribute Heads are responsible for matching the subject and the verb. We compute the dot product between the output $x$ of each attention head at the END token and $W_{U(v_{flip})}\!-\!W_{U(v)}$, the difference between the unembedding projections of the two verbs. Since the latter is fixed, we expect the projection of $x$ on it to be large, thus encouraging the probability difference between the two opposite verb forms. We visualize the dot production of the heads before and after fine-tuning in Figure 3. Through comparison, we observe an obvious flip at head.10.9, implying the reversal in its function from subject-verb agreement to disagreement. Other heads (3.0, 6.0, 4.4, 11.8, etc) also see flips with varying degrees, implying the self-adjustment of the functions inside the nodes. See Appendix D.4.1 for details.
Finding 3: The sharing of the Subject Identification Heads
The Subject Identification Heads attend heavily to the subject. One type of these heads attends to the END token (0.1, 0.3, etc), which is helpful to the cases like “they are”; the other type of heads (8.5, 10.5, etc) attends to the subject several tokens before, which is helpful to the cases like “the girl wearing a dress is”. We check the attention patterns in both types of these heads and find that they behave the same before and after fine-tuning. This implies that their functions are preserved and shared all the way down. This is consistent with the conclusion by Aghajanyan et al. (2020) that pre-training optimizes the description length without having direct access to the same tasks. The consistency further confirms the feasibility of our hypothesis in Section 3. In fact, these subgraphs share a significant portion of their structure (see Figure 11). These structures are crucial for the target task, while their function is shared before and after fine-tuning, thus they do not undergo the reversal seen in Finding 2.
Finding 4: The interaction and collaboration inside the model
Interestingly, we observe that over the training process, the nodes inside a circuit complete a task through a cooperative division of labor. Each head has its division of labor as discussed before. Besides, some heads directly affect the output (e.g. head.8.5), while others (e.g. head.7.4) cooperate with them to affect the output indirectly. Several heads achieve a common goal through cooperation: they could affect a head only through combined effect, and adjust themselves during training. These are found by knocking out a group of heads and check the change in the behaviors of others. See Appendix D.4.1 for details.
(a) circuit at 2000 steps
(b) circuit at 3000 steps
(c) circuit at 4000 steps
Figure 4: Phenomenon similar to Hebbian learning during the fine-tuning on the subject-verb disagreement task. We only present the Top-35 edges and the relevant nodes for clarity. The comparison among circuits at different steps helps us locate the key components for a task in a dynamic style.
Finding 5: Phenomenon akin to Hebbian learning
We visualize some of the circuits during fine-tuning in Figure 4. The thickness of an edge is proportional to its edge contribution. We observe a phenomenon akin to Hebbian learning (Do, 1949) that several edges are strengthened (edge contribution increased) during training, just as the synapses between neurons can be strengthened after continuous stimulation. Also, several edges are weakened at the same time (Table 3), which is similar to the lateral inhibition among neurons (Jacobson, 1993) that an activated neuron can reduce the activity of its neighbors. The change in an edge reflects the change in the direct effect from an upstream node to a downstream node. More details of this analogy are presented in Appendix D.4.3.
Through this analysis, we are able to locate the key components that the model really “cares about” in a dynamic style during fine-tuning. It is more rigorous than static circuit discovery methods when measuring the importance of a component, whereas the latter only focuses on the relative importance of the components in a model at a single, static state, while neglecting the trajectory of how this importance changes. It is this property that allows the model to develop a new circuit from the original “base circuit” (corresponding to the normal English grammar) during fine-tuning in a dynamic and heuristic way. Experiments in Section 5 further prove this.
4.5 Ablation studies and further discussions
The quality of the discovered circuit We first calculate the faithfulness and completeness Wang et al. (2022) of the circuits according to Equation 11 and 12. Faithfulness tells how much performance a circuit gets, while completeness tells how much performance a circuit fails to capture. As shown in Figure 8, the setting $N=1000$ reaches a relatively high faithfulness and extremely low completeness, demonstrating the effectiveness of our method to find the required parameters. We also observe an obvious turning point of the faithfulness and completeness at $N=1000$ as $N$ increases in Figure 8, which echoes the previous discussions in Section 4.3.
To further explore this, we plot the logit difference, PPL and the average tunable parameter ratio in Figure 5. The average tunable parameter ratio is computed during training. Every time after graph pruning, we recorded the number of the parameters $n_{i}$ corresponding to the nodes in the circuit, where $i$ is the count for graph pruning. After training, we calculated the average parameter ratio as $\frac{1}{M}\sum_{i=1}^{M}\frac{n_{i}}{N_{p}}$, where $M$ is the total number of the times for graph pruning, and $N_{p}$ is the total number of the parameters in the model. Similarly, we notice a common turning point of the three indicators at $N=1000$ in Figure 5. These findings all imply the shadow of the “intrinsic graph”.
@@ 4.5 Ablation studies and further discussions
Figure 5: The changes in indicators with various choices of NN (number of edges). The turning point at $N=1000$ implies the minimal of the subgraph for a target task at this point. Note that the parameter ratio varies and depends on the granularity of nodes. Theoretically more granular nodes result in lower parameter ratio because the location of the intrinsic nodes would be more precise.
Balancing the target task and general capabilities We provide an ablation study to randomly unfreeze the nodes outside the circuit during training and check if the performance is influenced. In practice, we randomly activate 10%, 20%, 30% and 40% of the parameters outside the final subgraph of subject-verb disagreement at $N=1000$, re-train the model and compare the results with before. As shown in Figure 9, we find that the performance is improved at the expense of harming other abilities, as the PPL on the validation set rises with the increase of irrelevant parameters. This implies that the “intrinsic graph” (the minimal subgraph) could strike a balance between the performance on the target task and the general capabilities. When a larger number of parameters are tuned, the target task could overwrite parameters that encode knowledge from other domains, thereby adversely affecting the general capabilities of the model.
5 Enhanced fine-tuning performance with circuit-tuning
5.1 Task descriptions and evaluation metrics
In this section, we test our method on larger models and more complex tasks. We apply circuit-tuning to Llama-3.2-1B/3B and Llama-3.1-8B (Dubey et al., 2024). We prepare two types of tasks based on whether reasoning is involved. For reasoning-based tasks, we focus on mathematics and logical reasoning. We use GSM8K (Cobbe et al., 2021) with zero-shot accuracy and Contexthub (Hua et al., 2024) with F1 score as datasets and metrics for mathematics and logical reasoning.
For reasoning-free tasks, only the final answer or a signal token is required. We prepare a gender de-biasing task which requires the language model to develop an unbiased perspective on genders, and a reading comprehension task which requires only the keywords as the answer. For the gender-debiasing task, we use BUG Levy et al. (2021) for training and WinoBias Zhao et al. (2018) for evaluation. We use the prejudice risk proposed in Liu et al. (2024) as the evaluation metric. For the reading comprehension task, we use SQuAD 2.0 Rajpurkar et al. (2018) with exact match and F1 score as evaluation metrics. Task settings and data examples are detailed in Appendix E.1.
To check if other capabilities are preserved during training, we evaluate on benchmarks involving general abilities as well as reasoning, coding, and multilingual abilities. For general abilities, we use MMLU Hendrycks et al. (2020), Winogrande Sakaguchi et al. (2021) and IFEval Zhou et al. (2023). For reasoning, coding and multilingual abilities, we use GPQA Rein et al. (2023), HumanEval Chen et al. (2021) and MGSM Shi et al. (2022) respectively. See Appendix E.3 for evaluation details.
5.2 Implementation details
Algorithm settings
For circuit-tuning, the graph settings of our method are the same as before in Section 4.2, except that each MLP layer is split into 64-dimensional “MLP heads” to achieve finer granularity. The settings of upstream / downstream nodes are presented in Table 8. As for the metric $\mathcal{L}_{m}$ for the quantification of a model’s prediction during circuit discovery, for gender de-biasing, we use the logit difference between male attribution words like $he/his$ and female attribution words like $she/her$. For other tasks, we simply set $\mathcal{L}_{m}$ as identical to the negative log probability loss for language modeling. Since the model abilities or behaviors for completing these task are hidden among multiple tokens, we focus on the overall ability in terms of a task instead of manually separating out a single capability (e.g. the Add capability in math). For gender de-biasing, We selectively add $\beta\cdot|\mathcal{L}_{m}|$ that acts as a regularization term in the loss $\mathcal{L}$ for explicit guidance.
Choice of hyper-parameters
Previous discussions in Section 4.5 point out that circuit-tuning has the potential to preserve general capabilities during fine-tuning. To further verify this, we compare it with full fine-tuning and LoRA. For LoRA, we sweep over a wide range of values and set rank $r=32$ and $\alpha=64$ for all experiments, since the task performance is the best under this setting. For circuit-tuning, we set $K=8$, i.e., we perform circuit discovery every 8 steps of optimization for efficiency. As for the decision of the number of edges $K$ in a circuit, we perform graph pruning under several choices of $N$ on a batch of samples before fine-tuning. Then we calculate the faithfulness and completeness under each setting. This will result in a curve with a knee-point similar to that in Figure 8. The value of $N$ at the knee-point is what we use. In pratice, we use 2000, 3000, and 4000 for 1B/3B/8B models, respectively. Details for other hyper-parameters are shown in Appendix E.2.
5.3 Main results and discussions
Table 1: Evaluation on complex tasks. All results are the average of 10 runs with random sampling.
Methods & Tasks Math Logical Reasoning Gender De-biasing Reading General Capabilities Computation
@@ 5.3 Main results and discussions
Acc@1 (%) F1 (%) Prejudice Risk $\downarrow$ Exact Match(%) Performance Change (%) Avg Param Ratio
Llama-3.2-1B-it 40.71 20.35 0.555 39.68 / /
Llama-3.2-1B-it-full-tuning 46.47 26.89 0.533 36.73 0.14 1.00
Llama-3.2-1B-it-lora 44.58 22.51 0.530 34.30 -8.55 1.79e-2
Llama-3.2-1B-it-circuit-tuning 45.56 27.06 0.312 41.78 2.33 7.65e-2
Llama-3.2-3B-it 70.36 42.71 0.641 54.12 / /
Llama-3.2-3B-it-full-tuning 75.44 47.32 0.632 55.63 -2.54 1.00
Llama-3.2-3B-it-lora 73.54 46.27 0.638 54.73 -3.92 1.49e-2
Llama-3.2-3B-it-circuit-tuning 74.35 47.59 0.417 56.58 -1.80 8.57e-2
@@
Llama-3.1-8B-it-full-tuning 83.76 49.53 0.640 59.60 -2.76 1.00
Llama-3.1-8B-it-lora 80.21 47.64 0.643 58.97 -4.15 1.03e-2
Llama-3.1-8B-it-circuit-tuning 82.97 50.54 0.420 60.04 -0.94 9.37e-2
Table 1 presents the main results of the experiment. For fairness, all results are the average of 10 runs with random sampling. The “performance change” is calculated as the rate of change in performance of the models after fine-tuning compared to those before fine-tuning (averaged across all benchmarks). It can be observed that circuit-tuning achieves strong performance on all four tasks, and obtains the best performance on three of them except math. For math, the performance of circuit-tuning is superior to that of LoRA but does not match full fine-tuning. This reflects a key property of circuit-tuning: it strikes a trade-off between target tasks and general capabilities. Compared to other tasks, the abilities required for math are more specialized, and the math expressions differ significantly from natural language. Therefore, when we select the $N$ corresponding to the knee-point to limit the subgraph size in circuit-tuning, the model will constrain parameter updates as much as possible to the part of the computational graph responsible for math, while minimizing the impact on other parts. This avoids improving target task performance by co-opting parameters from other capabilities, which would otherwise harm general capabilities. This point is also supported by the general capability results in Table 1, which show that our method is better at preserving them.
In Table 1, we also report the average tunable parameter ratio, the calculation of which is detailed in Section 4.5. It is evident that, compared to full fine-tuning, our method requires updating a much smaller number of parameters, focusing more on task-related parts. Although circuit-tuning requires more computation than LoRA, it achieves better performance on both the target task and general capabilities. Note that when we increase the number of parameters for LoRA (by increasing the rank and adjusting $\alpha$), its performance does not improve. This reflects a limitation of LoRA: although it updates parameters in a subspace, it does not explicitly select parameters relevant to the target task. In contrast, circuit-tuning can adaptively select the parameters during fine-tuning. For instance, if certain parameters receive more adjustments in the early stages of training but need to remain stable later on, the corresponding nodes in the computational graph will then be discarded.
Figure 6: The prejudice risk before and after circuit-tuning. A regularization term in the loss could help to modify a model’s stereotype.
Through the experiments in this section, we have demonstrated that circuit-tuning can be extended to larger models and more complex tasks. From an interpretability perspective, we can analyze the learning dynamics using methods similar to those in Section 4. For math tasks, for example, we can analyze the changes in edge contribution with the number of training steps to locate the key parameters responsible for the task. This allows for flexible interventions and more precise fine-tuning for that task. Similarly, as described in Section 5.2, we selectively added a regularization term related to the target capability to the loss function for the gender de-biasing task, as shown in Equation 15. The prejudice risks of the models before and after gender de-biasing are visualized in Figure 6. Following Liu et al. (2024), we regard the distribution of prejudice risk as a normal distribution over the 40 types of occupations in the WinoBias test set and perform interpolation on the computed results. The dynamic process of de-biasing can be observed from right to left in Figure 6. It is obvious that with a regularization term in the loss function, the distribution of the prejudice risk is more concentrated to a smaller value. Thus we can customize the algorithm settings in circuit-tuning flexibly according to the requirement of a task, demonstrating its flexibility and effectiveness in practice.